OpenStudy (prince1342):

A textbook store sold a combined total of 467 sociology and physics textbooks in a week. The number of physics textbooks sold was 41 less than the number of sociology textbooks sold. How many textbooks of each type were sold?

MARC (marc_d):

Total textbook sold=467 physics textbooks sold=\[\frac{ 467-41 }{ 2 }\]

MARC (marc_d):

\[=\frac{ 426 }{ 2 }\]\[=213\]

MARC (marc_d):

Sociology textbooks sold=467-213 =?

MARC (marc_d):

solve it @Prince1342

OpenStudy (danjs):

The information given gives you a couple of equations, the total number of books sold, and, number of physics compared to number sociology books sold. The total number sold is the combined total of physics and sociology sold #total books sold = #physics sold + #sociology sold They relate physics sold to sociology sold from the second statement. The number of physics textbooks sold was 41 less than the number of sociology textbooks sold. #Physics sold = #sociology sold - 41

OpenStudy (danjs):

To make it simpler, let Physics books sold be P and Sociology sold be S, the two equations from the problem are 467 = P + S and P = S - 41 ------------- you can use the second equation and substitute what S is into the first equation. 467 = P + S 467 = S - 41 + S 467 = 2*S - 41

OpenStudy (danjs):

Solving you get S=254, and so using that value, you get 467 = P + S 467 = P + 254 P = 213

MARC (marc_d):

*thumbs up* @DanJS

OpenStudy (prince1342):

Thank you!